Consider the Harrod-Domar model given by the following equations.

S t = 0.3 Y t

I t = 0.25 (Y t - Y t - 1)

I t = S t


a) Explain the significance of each variable and of all the constants. 

b) Use the same constants to write the continuous-time Harrod-Domar model. 

c) If Y 0 = 500, derive the income trajectory for the given discrete model.

d) Compute the fixed point of the trajectory determined above and analyze the system’s stability



1
Expert's answer
2021-05-30T14:15:15-0400

a) St- saving in t period; It- investment in t period; Yt/t-1- GDP/economy outcome in t period/previous period.

b) "0.3Y_{t-1}=0.25(Y_t-Y_{t-1})"

"Y_t=2.2Y_{t-1}"

c) "Y_1=2.2*500=1100"; "Y_2=2.2*1100=2420" it is exponential growing function.

d) "Y_0=500" is fixed point. The system is not stable because the function move away from the fixed point.